OpenStudy (anonymous):

arush sat down to study at 8 a.m. the angle of elevation of the sun was 15'.when he finished studying the angle of elevation of the sun was 55'.how long did arush study?

OpenStudy (anonymous):

@Hero @karatechopper

OpenStudy (anonymous):

plz hlp me out today is my xam PSA

OpenStudy (anonymous):

@sirm3d @onedirection822 @Weepdo-s @poopsiedoodle

OpenStudy (anonymous):

timo u r writing something were that

OpenStudy (anonymous):



OpenStudy (anonymous):

be nice poopsie how the heck you become ambassador?

OpenStudy (anonymous):

set a ratio original to new time 8/x=15/55

OpenStudy (anonymous):

that reduces to 8/x = 3/11

OpenStudy (anonymous):

or i think i did it wrong

OpenStudy (anonymous):

is there more info ruchi?

OpenStudy (anonymous):

yes option is also given 40 mins, 1hrs 20mins, 2hrs 23mins, 2hrs 40 mins

OpenStudy (anonymous):



OpenStudy (anonymous):

ruchi there should be a more info. other than choices.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i am confused perhabs if we divide the angle 360 into a 24 hour thingy. so 360 degree= 24 hour

OpenStudy (anonymous):

nt getting